OpenStudy (anonymous):

An integer from 1 to 40 is drawn at random.Find the probability of each result. odd AND multiple of 5 _____ odd OR a multiple of 5______ Please show me the work

OpenStudy (anonymous):

1) is it 4/40 =1/10 2) 24/40 = 3/5

OpenStudy (anonymous):

how many odd numbers are there? how many multiples of five are there? 1. (# of odd numbers that are multiples of 5)/total amount of numbers 2. (# of odd numbers + # of multiples of 5)/total amount of numbers

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

but for 2) (no. of odd numbers + no of multiples of 5 - no ofodd multiples of 5)/total no. of numbers

OpenStudy (anonymous):

Thanks matricked ! I got it.